11. Write the standard equation of the circle with center (-3, 5) and a diameter of 18.
I am begging anyone for help with this please !

Answer by htmentor(1343)

The equation of a circle with center (a,b) and radius R is
(x-a)^2 + (y-b)^2 = R^2
We are given the center and the diameter, which is 2R
So R = 9 and the equation is (x+3)^2 + (y-5)^2 = 81